The supply function of a certain product is P = 16 + 2Q, where P is the price and Q is the number of units produced. Find the producer surplus if the market price P = 200. [1] 7 400 [2] 8 464 [3] 12 464 [4] 20 000


1
Expert's answer
2021-10-01T09:49:06-0400

Producer surplus = "1/2\cdot Q(P_1-P_2)"

where Q is quantity, P1 is price,

P2 is minimum price any producer is willing to sell at.


"Q=\frac{P-16}{2}=\frac{200-16}{2}=92"

"P_1=200"

"P_2=16"

Producer surplus ="1/2\cdot 92\cdot(200-16)=8464"

Answer: [2] 8 464